# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Results
Together, the analysis of the null models sheds light on the factors underlying gender differences in the acquisition and diffusion of knowledge over time. Specifically, we find that unfairness is not driven by the gap in the number of women, rather it arises from the structure of the collaboration network.
@@ Discussion
In this work, we investigated gender inequalities in the dissemination of scientific knowledge through the APS collaboration network. Although, in principle, knowledge flow is not contingent on gender, we discovered a systematic trend suggesting that women may face challenges in acquiring and diffusing knowledge. These two dimensions of unfairness are not symmetric: Whereas the difference in knowledge diffusion is less marked and seems to improve over time, women’s disadvantage in knowledge acquisition is considerable and persistent. This disparity is critical, as it might prevent women from accessing novel ideas, new collaborations, or even finding job opportunities. Exploring the potential drivers of unfairness, we found that the gap in the number of female scholars does not explain women’s disadvantage; instead, the difference in the cumulative number of collaborators jeopardizes women’s opportunities to acquire and diffuse knowledge.
@@
These findings challenge the hypothesis that increasing the number of women in research is enough to provide equal opportunities to acquire and diffuse scientific knowledge. This lack of opportunities is not simply related to when women enter the system but to how they take part in it, that is, the evolution of their network of collaborators. Differences between women and men in the size of their networks may stem from multiple factors. First, as a historically disadvantaged minority in a male-dominated field and with more parental responsibilities, women might be excluded from cutting-edge collaboration opportunities [42, 43]. Second, women are more likely to drop out of academia in all career ages compared to men [4], preventing them from growing their network (see Fig. S4 in Supplementary Information for the analysis of career lengths in our data). Moreover, women tend to have less academic mobility than men [44]: Since mobility plays a crucial role in expanding scholars’ network, this might contribute to widen the gap in the number of collaborators. Remarkably, these differences are intertwined with time, which results in slowing down the acquisition and diffusion of knowledge, thus creating a temporal glass ceiling [45] for women in physics.
@@
Our work sheds light on how the dynamics of scientific collaborations shape gender disparities in knowledge dissemination and calls for a deeper understanding on how to intervene to guarantee fairness and diversity in the scientific community. Specifically, it highlights that policies intended to increasing the number of female scholars in academia may not be sufficient to reduce gender disparities in the acquisition and diffusion of scientific and technical knowledge; instead, it points to interventions towards prevention of women career dropout and promotion of academic mobility and diverse collaborations.
Methods
@@
We construct the collaboration network from 678,916 publication records of the American Physical Society, ranging from $1893$ to $2020$. This dataset contains, for each publication, a list of coauthor names and their affiliations, alongside other metadata. We apply a recently developed implementation [36] of an existing name disambiguation solution [35] to link publications to authors. For a pair of publications with matching author names, this solution uses affiliations, citations, and coauthors to decide whether the two names and the respective publications belong to the same author. It applies a conservative strategy by merging only authors for which there are no conflicting first names. This reduces the number of authors with extreme publication counts. In the collaboration network, this limits the number of nodes with very high degree, which could deeply affect the contagion dynamics.
Similar to previous studies [11, 36, 4], we infer gender labels from author names. Using a recently developed open-source solution [37], we estimate 34,596 scientists to be women, 195,490 to be men, while the gender of 200,669 disambiguated authors remains unknown. For authors with multiple names, we remove names for which no gender could be inferred and assign gender based on a majority vote, i.e., if the majority of names are labeled as male (female), the scientist is considered male (female). Authors with a draw in the number of female and male labeled names or those for whom only “unknown” labels could be inferred, most often due to providing only initials or unisex names, are labeled as “unknown”. Note that these authors remain present in the network, namely they participate in the knowledge spreading process. However, they are ignored when 1) seeding the contagion process and 2) evaluating the fairness of the system.
@@
We explore the factors underlying the emergence of acquisition and diffusion unfairness in the flow of knowledge by designing three null models of the APS collaboration network. For each model, the structure of the temporal network, namely the nodes and the connections between them in every snapshot, is unchanged while the node attributes, meaning the authors’ gender, are reshuffled so as to preserve certain characteristics of the system.
  • Balance. This model generates an ideal scenario where 1) 50% of the authors are women at any time; 2) men and women do not differ in any network feature. Starting from the first snapshot of the temporal network, we randomly assign the gender to the nodes with a 50/50 ratio. In the following snapshots, old nodes keep their gender, while the gender of new nodes is again attributed randomly with a 50/50 ratio. The main purpose of this model is to validate our measures of unfairness, as the scenario it creates should be perfectly fair.
  • Late arrival. This model preserves the proportion of women in each year, so as it remains the same as in the data, and removes gender differences in the network features. We reshuffle the gender of the nodes in the first snapshot of the network. In the following snapshots, we keep the gender of old nodes and reshuffle that of new ones. The model allows us to investigate how the late arrival of women in the collaboration network affects their opportunity to acquire and diffuse scientific knowledge, while controlling for the structure of the network.
  • Coauthors. With this model, we preserve the proportion of women in each year and the difference between genders in the distribution of the number of coauthors. First, we order nodes by their total number of connections and divide them into groups, each representing 2% of the nodes. As for the Late arrival model, we start from the first snapshot and reshuffle gender of nodes within the same group. In this way, we explore how gender difference in the distribution of coauthors impacts the fairness of the system on top of the late arrival of women in the network.
Data Availability
The APS datasets are available upon request to the American Physical Society (https://journals.aps.org/datasets).
@@ Acknowledgments
C.Z. acknowledges the support of 101086712-LearnData-HORIZON-WIDERA-2022-TALENTS-01 financed by EUROPEAN RESEARCH EXECUTIVE AGENCY (REA) (https://cordis.europa.eu/project/id/101086712). C.Z. and L.G. thank César Hidalgo and Balázs Lengyel for the insightful comments and discussion. L.G. and F.B. acknowledge support of the Air Force Office of Scientific Research under award number FA8655-22-1-7025. J.B. was supported by the Austrian Science Promotion Agency FFG under project No. 873927 ESSENCSE. and is a recipient of a DOC Fellowship of the Austrian Academy of Sciences at the Complexity Science Hub.

### v2

@@ Results
Together, the analysis of the null models sheds light on the factors underlying gender differences in the acquisition and diffusion of knowledge over time. Specifically, we find that unfairness is not driven by the gap in the number of women, rather it arises from the structure of the collaboration network that creates a cumulative disadvantage for women over time.
@@ Discussion
In this work, we investigated gender disparities in the dissemination of scientific knowledge through the APS collaboration network. Although, in principle, knowledge flow is not contingent on gender, we discovered a systematic trend suggesting that women may face challenges in acquiring and diffusing knowledge. These disparities are critical: On the one hand, late knowledge acquisition can generate a competitive disadvantage that may prevent women from accessing novel ideas, new collaborations, or even finding employment; on the other hand, slower knowledge diffusion may negatively affect women’s visibility, fostering gender stereotypes and impacting their academic recognition, exacerbating the Matilda effect  (?). Our findings suggest that these two dimensions of unfairness are not symmetric: Whereas the difference in knowledge diffusion is less marked and seems to improve over time, women’s disadvantage in knowledge acquisition is considerable and persistent.
Exploring the potential drivers of unfairness, we found that the gap in the number of female scholars does not explain women’s disadvantage; instead, the difference in the cumulative number of collaborators jeopardizes women’s chances to acquire and diffuse knowledge. Beyond knowledge dissemination, different collaboration patterns could limit opportunities for women throughout their careers, as job and promotion offers are often the result of peer-to-peer interactions.
@@
These findings challenge the hypothesis that increasing the number of women in STEM is enough to provide equal opportunities to acquire and diffuse scientific knowledge. This lack of opportunities is not simply related to when women enter the system but to how they take part in it, that is, the evolution of their network of collaborators. Differences between women and men in the size of their networks may stem from multiple factors. First, as a historically disadvantaged minority in a male-dominated field and with more parental responsibilities, women might be excluded from cutting-edge collaboration opportunities  (?, ?, ?). Second, women are more likely to drop out of academia in all career ages compared to men  (?), preventing them from growing their network (see Fig. S4 in Supplementary Information for the analysis of career lengths in our data). Moreover, women tend to have less academic mobility than men  (?): Since mobility plays a crucial role in expanding scholars’ network, this might contribute to widening the gap in the number of collaborators. Remarkably, these differences are intertwined with time, which results in slowing down the acquisition and diffusion of knowledge, thus creating a temporal glass ceiling  (?) for women in physics.
@@
Our work sheds light on how the dynamics of scientific collaborations shape gender disparities in knowledge dissemination and calls for a deeper understanding on how to intervene to guarantee fairness in the scientific community. Specifically, it highlights that policies intended to increase the number of female scholars in academia may not be sufficient to reduce gender disparities in the acquisition and diffusion of scientific and technical knowledge. In line with the policy changes recently suggested by the scientific community  (?, ?, ?, ?), our findings point to institutional interventions that prevent women’s career dropout and encourage academic mobility and diverse collaborations, with the aim of promoting equity, diversity, and inclusion in science.
Materials and Methods
@@
We construct the collaboration network from 678,916 publication records of the American Physical Society, ranging from $1893$ to $2020$. This dataset contains, for each publication, a list of coauthor names and their affiliations, alongside other metadata. We apply a recently developed implementation  (?) of an existing name disambiguation solution  (?) to link publications to authors. For a pair of publications with matching author names, this solution uses affiliations, citations, and coauthors to decide whether the two names and the respective publications belong to the same author. It applies a conservative strategy by merging only authors for which there are no conflicting first names. This reduces the number of authors with extreme publication counts. In the collaboration network, this limits the number of nodes with very high degree, which could deeply affect the contagion dynamics.
Similar to previous studies  (?, ?, ?), we infer gender labels from author names. Using a recently developed open-source solution  (?), we estimate 34,596 scientists to be women, 195,490 to be men, while the gender of 200,669 disambiguated authors remains unknown. For authors with multiple names, we remove names for which no gender could be inferred and assign gender based on a majority vote, i.e., if the majority of names are labeled as male (female), the scientist is considered male (female). Authors with a draw in the number of female and male labeled names or those for whom only “unknown” labels could be inferred, most often due to providing only initials or unisex names, are labeled as “unknown”. Note that these authors remain present in the network, namely they participate in the knowledge spreading process. However, they are ignored when 1) seeding the contagion process and 2) evaluating the fairness of the system.
@@
We explore the factors underlying the emergence of acquisition and diffusion unfairness in the flow of knowledge by designing three null models of the APS collaboration network. For each model, the structure of the temporal network, namely the nodes and the connections between them in every snapshot, is unchanged. In contrast, the node attributes, meaning the authors’ gender, are reshuffled to preserve certain characteristics of the system.
  • Balance. This model generates an ideal scenario where 1) 50% of the authors are women at any time; 2) men and women do not differ in any network feature. Starting from the first snapshot of the temporal network, we randomly assign the gender to the nodes with a 50/50 ratio. In the following snapshots, old nodes keep their gender, while the gender of new nodes is again assigned at random with a 50/50 ratio. The main purpose of this model is to validate our measures of fairness, as the scenario it creates should be perfectly fair.
  • Late arrival. This model preserves the proportion of women in each year so that it remains as it is in the data and removes gender differences in the network features. We reshuffle the gender of the nodes in the first snapshot of the network. In the following snapshots, we keep the gender of old nodes and reshuffle that of new ones. The model allows us to investigate how the late arrival of women in the collaboration network affects their opportunity to acquire and diffuse scientific knowledge, while controlling for the structure of the network.
  • Coauthors. With this model, we preserve the proportion of women in each year and the difference between genders in the distribution of the number of coauthors. First, we order nodes by their total number of connections and divide them into groups, each representing 2% of the nodes. Analogously to the Late arrival model, we start from the first snapshot and reshuffle the gender of nodes within the same group. In this way, we explore how gender differences in the distribution of coauthors impact the fairness of the system on top of the late arrival of women in the network.
@@ Acknowledgments
C.Z. acknowledges funding by the European Union under Horizon EU project LearnData, 101086712. C.Z. and L.G. thank César Hidalgo, Balázs Lengyel, Vito Latora and Esteban Muñoz for the insightful comments and discussion. L.G. and F.B. acknowledge support of the Air Force Office of Scientific Research under award number FA8655-22-1-7025. J.B. was supported by the Austrian Science Promotion Agency FFG under project No. 873927 ESSENCSE and the DOC Fellowship of the Austrian Academy of Sciences.
Competing interests
The authors declare they have no competing interests.
Data and material availability
The APS datasets are available upon request to the American Physical Society
(https://journals.aps.org/datasets).